OpenStudy (anonymous):

please help :( If f(x) = −2x − 4 and g(x) = x + 4, find (f/g)(−8).

OpenStudy (anonymous):

well to me it looks like: -2x-4/x+4 =-2(-8)-4/-8+4 =16-4/-4 =-3

OpenStudy (anonymous):

Thanks :)